# Structural response of neutron stars to rapid rotation and its impact on the braking index

### v1

@@ Abstract
Pulsars are rotating neutron stars that are observed to be slowing down, implying a loss of their kinetic energy. There can be several different physical mechanisms involved in their spin-down process. The properties of fast-rotating pulsars depend on the nature of the neutron star matter, which can also affect the spin-down mechanisms. In this work, we examine three different physical phenomena contributing to the spin-down: magnetic dipole radiation, gravitational mass quadrupole radiation due to the “mountain” formation, gravitational mass current quadrupole radiation or the r-modes, and calculate the expressions for the braking indices due to all of them. We have also considered jointly the implications of the uncertainties of the equation of the state of neutron star matter and rapid rotation on the braking indices corresponding to the aforementioned processes and their combinations. In all cases, the rapid rotation results in a departure from the standard values in the literature for the braking index when the rotational effects are ignored. If generated with a saturation amplitude within the range of $10^{-4}-10^{-1}$, the r-mode oscillations dominate the spin-down of millisecond pulsars. We also explore the braking index in the context of millisecond magnetars. This study examines two braking index measurements in the context of newly born millisecond magnetars from two observed short $\gamma$-ray bursts. The measured braking indices for these objects are consistent with our estimation, which allows us to conclude that the spin frequency of the remnants is within the range of 550850similar-toabsent550850$\sim$ 550–850∼ 550 – 850 Hz.
@@ I Introduction
The majority of neutron stars (NS) are observed as pulsars, which are rapidly rotating and observed to spin down over time, indicating a loss of rotational kinetic energy. The spin-down rate is given as $\dot{\Omega}=-k\Omega^{n}$ Manchester et al. (1985), where $n$ is the braking index and $k$ is a positive constant. The braking index captures information about the mechanism of energy loss. Depending on the emission loss mechanism $k$ can be a function of radius, moment of inertia, magnetic field strength etc (discussed later in the text). In most cases, the pulsar’s rotational energy loss is attributed to the magnetic dipolar radiation (MDR) from the pulsar magnetosphere for which the braking index is 3 Gunn and Ostriker (1969). However, pulsars can lose energy via gravitational wave (GW) radiation if there is non-axisymmetric deformation or r-mode instability, for which the braking index is 5 and 7 respectively (Riles, 2023). Therefore, it is natural to expect a rapidly rotating neutron star to lose energy via all three channels if the conditions are favourable. Apart from these three channels of energy loss, pulsars also lose energy due to the particle wind (Harding et al., 1999; Spitkovsky, 2006).
@@
The spin-period of pulsars spans a large range. The slowest pulsar has a rotation period of 76 s Caleb et al. (2022) to the most rapidly rotating being 1.39 ms Hessels et al. (2006). The total population is typically categorized as normal (or “slow”) pulsars and millisecond pulsars, with the demarcation around 16 ms Halder et al. (2023). The braking index has been measured for a few younger pulsars and found be less than 3 Livingstone et al. (2007); Weltevrede et al. (2011); Livingstone and Kaspi (2011); Lyne et al. (1993); Boyd et al. (1995); Roy et al. (2012); Espinoza et al. (2011), which is often attributed to the evolution of the magnetic field, modified magnetosphere or even the interaction of supernova fallback disk Chen and Ruderman (1993); Melatos (1997); Harding et al. (1999). A similar sub-3 braking index can be achieved in rapidly rotating pulsars, which is purely due to structural change in the star due to rapid rotation. In such a scenario, the $k$ cannot be assumed independent of the spin frequency Hamil et al. (2015), which has been discussed in detail in Sec. IV.
@@
The effect of structural evolution is prominent over a spin frequency of 200 Hz Hamil et al. (2015). Currently, the ATNF Pulsar Catalogue  [^1] https://www.atnf.csiro.au/research/pulsar/psrcat/ Manchester et al. (2005) records 392 pulsars with spin-frequency more than 200 Hz, which corresponds to $\sim 11\%$ of the total pulsar population discovered till date. These rapidly rotating pulsars are important in probing the physics of dense nuclear matter. The physics of the nuclear interaction and the constituents of the matter determines the dense matter equation of state (EOS), which in turn governs the stellar structure and its response to rapid rotation. To study the frequency evolution of the stellar structure and the braking index, we have adopted a semi-agnostic approach to construct the dense matter equation of state following the formalism of Gandolfi et al. (2019), which is summarized in Sec. II.
@@
There are various types of constraints that can be applied to the EOS model. Astrophysical observations provide several constraints relevant to the high density part of the EOS. In particular, the radio observations of the most massive pulsar, PSR J0740+6620 Cromartie et al. (2019); Fonseca et al. (2021) provide the most stringent constraint on the NS EOS. The tidal deformability measurements from the multimessenger observations of the binary neutron star merger event GW170817 put additional constraints on the EOS Abbott et al. (2017a, b, 2019a). In recent years, the Neutron star Interior Composition Explorer (NICER) collaboration has reported a few simultaneous mass-radius measurements of PSR J0030+0451, PSR J0740+6620, and PSR J0437-4715, respectively, that also help us to constraint the Mass-radius plane and consequently the EOS parameter space Miller et al. (2019); Riley et al. (2019); Miller et al. (2021); Riley et al. (2021); Salmi et al. (2022); Vinciguerra et al. (2024); Choudhury et al. (2024). On the other hand, the latest constraint on the low-density part of the EOS comes from the advancements of the chiral effective field theory ($\chi$-EFT) calculations Tews et al. (2013); Hebeler et al. (2013); Lynn et al. (2016); Drischler et al. (2019); Huth et al. (2021).
@@
Figure 2: The brown lines indicate the constrained EOSs ($\sim$6K) corresponding to the posterior distributions of figure 1 and the blue lines indicate the EOS which falls within the 68% confidence interval of the constrained posterior distribution of the EOS parameters.
Figure 3: The M-R diagrams for all $\sim$6K EOSs from figure2 have been shown in the brown coloured lines and the blue lines indicate the M-R relations corresponding to the 68% confidence interval of the posterior distribution of the constrained EOS parameters.
@@ II Description of neutron star matter
where, $P(m_{1},m_{2}|\mathbf{X})$ is the prior distribution for the component masses of the binary. For simplicity, we have chosen a uniform prior for $m_{1}$ and $m_{2}$. We use a Gaussian kernel density estimator to construct the GW likelihood from the discreet data. We have fixed the chirp mass to its median value because of the high precision of the measurement. Then, we construct binaries corresponding to that chirp mass by varying $m_{1}$ and determine the corresponding $m_{2}$. For each EOS, we compute the tidal deformabilities ($\Lambda_{1},\Lambda_{2}$) for the pairs of ($m_{1},m_{2}$), and find the probability using equation II.
@@
Finally, we sample the posterior using a nested sampling algorithm Skilling (2006), in the dynesty software package Speagle (2020). The posterior distributions of the EOS parameters for nuclear matter are shown in Fig. 1. We have also included the derived values for $E_{\rm sym}$ and $L$ at $\rho_{0}$. We have found $E_{\rm sym}=31.49^{+3.43}_{-3.11}$ MeV and $L=58.82^{+10.24}_{-8.44}$ MeV at 68% CI. These values are well within the agreed ranges found in literature Oertel et al. (2017); Margueron et al. (2018). This affirms the robustness of our EOS samples used to study the rotational properties of pulsars in this work. We have shown the corresponding EOSs and their mass-radius curves in Fig. 2 and 3, respectively. We can see from Fig. 3 that the EOSs generated in our exercise span a large range of radii for a $1.4M_{\odot}$ star. Therefore, these curves are able to represent the parameter space of neutron star matter. Note that, we have not used the data from NICER measurements in our analysis as we see that our 68% mass-radius curves in Fig. 3 are consistent with 68% of those sources.
@@ III Rotational effects on the structure of Neutron stars
The rapid rotation generates a large amount of centrifugal force, which leads to the departure from the spherical shape of the star that is characterised by the ratio of the polar to the equatorial radius $R_{p}/R_{e}$. The dependence of this ratio on frequency is shown in the bottom left panel of Fig. 4 for all EOS.lowest frequencies for which the RNS code converges. Subsequently, we generate the Due to the deformation in the shape of the star, the definition of the radius becomes ambiguous in such a case, yet, in many scenarios, for estimating various physical quantities, the radius of the star is required as discussed in Sec. IV. Hence, in this work, we define a frequency-dependent radius by averaging over the circumferential radius of a rotating star given as
@@
The expression of circumferential radius $R(\theta,\Omega)=R_{e}[1-\bar{\Omega}(0.788-1.03x)\cos^{2}\theta]$ is taken from the previous work by AlGendy and Morsink (2014) and Suleimanov et al. (2020), where $\theta$ is the polar angle, $\bar{\Omega}=\Omega(R^{3}_{e}/GM)^{1/2}$ and $x=GM/c^{2}R_{e}$. The EOS and rotation frequency-dependent average radius of the star is shown in the lower right panel of Fig. 4, which can be seen to decrease strongly at higher frequencies ($\gtrsim 850$Hz). Also at the higher frequency, the $R_{p}$ can be $\sim 25\%$ smaller than the $R_{e}$, therefore our method of computing a unique value of a radius and attributing a spherical shape to the star may have some shortcomings even if it mathematically renders a valid solution. In the upper left and right panels of Fig. 4 we show the variation of the gravitational mass and moment of inertia as a function of the rotation frequency for all EOS. It is evident from Fig. 4 that all these physical quantities have significant frequency dependence and will contribute to the braking index of the NS which we will discuss in the next section. We have confined our studies to the rotational frequency above $\gtrsim 200$ Hz following Hamil et al. (2015), where the frequency dependence of the macroscopic quantities has been shown to be prominent above $\gtrsim 200$ Hz only.
@@
The high-frequency behaviour of the macroscopic quantities observed in Fig. 4 is controlled by the stiffness of the EOSs. At a given density, a stiffer EOS produces more pressure than a softer one and, therefore, can defy gravity more easily. Hence, if there is no rotation, the stiffer EOS will make a star of a larger radius and a lower central density if a certain number of baryons (equivalent to a fixed baryon mass) are put together. Consequently, the star with the stiffer EOS will be less bound, resulting in a higher gravitational mass. Note the gravitational binding energy is defined as $M_{b}-M$. This behaviour remains even if the stars rotate. Therefore, the upper curves in $M$-vs-$f$, $R$-vs-$f$, and $I$-vs-$f$ plots in Fig. 4 correspond to stiffer EOSs. With increasing frequency, the equatorial radii of all the stars (irrespective of the underlying EOS) grow, making them less bound and having higher gravitational mass. Furthermore, since at a given rotational frequency, the star with the stiffer EOS is less bound than the softer EOS, it becomes unstable at a lower frequency. In other words, the Kepler frequency is smaller in the case of stiffer EOSs. It explains the behaviour seen in the $M-f$ plot. The $R_{p}/R_{e}$-vs-$f$ curve captures the evolution of the stellar deformation as the function of rotation frequency. It is clear from the above discussion that for the stiffer EOSs, the equatorial radius increases much faster with the small change to the rotation frequency compared to the softer EOSs. Consequently, the ratio $R_{p}/R_{e}$ falls faster in the case of stiffer EOSs. Therefore, the left curves in the $R_{p}/R_{e}$-vs-$f$ plot belong to stiffer EOSs.
@@ IV.1 Calculation of braking index
$|\dot{E}_{\rm{EM}}|=\frac{2}{3}\mu^{2}\Omega^{4}\sin^{2}\alpha\equiv\mathcal{E}_{EM}R^{6}\Omega^{4},$ (12)
where the magnetic moment $|\mu|^{2}=B_{p}^{2}R^{6}/4$, with the surface magnetic field strength $B_{p}$, stellar radius $R$, $\alpha$ is the angle of inclination between the rotation and the magnetic axis and $\mathcal{E}_{EM}=B_{p}^{2}\sin^{2}\alpha/6$. The corresponding spin-down rate using Eq. 10 leads to
$\dot{\Omega}_{\rm MDR}=-\frac{2B_{p}^{2}R^{6}\sin^{2}\alpha}{3I}\Omega^{3}\Rightarrow-k_{\rm MDR}\Omega^{3},$ (13)
indicating the MDR leads to braking index 3. However, in Sec. III we have noticed at higher spin-frequency the $R$ and depends on $f$ making $k_{\rm MDR}$ dependent on $f$ and hence the braking index. On the other hand, the $f$ dependence of $I$, will not allow us to arrive at the expression 13, as $dI/d\Omega\neq 0$. Hence, for a rapidly rotating star, the modification to the spin-down expression is essential.
@@
The expression of $n(\Omega)$ above shows the evolution of the braking index with the spin-frequency, which is dependent on the EOS through the frequency derivative of the macroscopic quantities like the $M$, $R$ and $I$. The $n_{\rm EM}$ term is the contribution of MDR loss, $n^{Q}_{\rm{GW}}$ is the contribution from the GW radiation due to long-lived deformations, $n^{R}_{\rm{GW}}$ is the contribution from the emission loss via the r-mode oscillations. The term $n_{I}$ depends purely on the moment of the inertia and its higher order derivatives, which has been previously derived in Glendenning et al. (1997). The quantities $\mathcal{E}_{EM},\epsilon$ and $\alpha_{\rm S}$ can be considered as the switches which by setting to zero help in analysing the braking index in the absence of a particular channel of energy loss. For example, setting $\alpha_{\rm S}=0$ will help in understanding the evolution of the braking index purely due to the electric dipolar emission loss and the GW emission loss due to finite $\epsilon$. In the limit of slow rotation where all the derivatives w.r.t $\Omega$ tend to zero, by setting $\epsilon$ and $\alpha_{\rm S}$ to zero we obtain $n=3$, by setting $\mathcal{E}_{EM}$ and $\alpha_{\rm S}=0$, we obtain braking index 5 and finally by setting $\mathcal{E}_{EM}=\epsilon=0$, we obtain $n=7$, which are in agreement with the limiting cases known in the literature.
Figure 5: This figure shows the effect of rotation on the braking index of pulsars (see Sec. III for detailed discussions and references). The left panel indicates the braking index purely due to magnetic dipolar radiation of magnetic field strength $B_{p}=10^{8}$ G and $\alpha=45^{\circ}$. The middle panel shows the variation of the braking index if the energy loss happens only through the GW emission in the presence of long-lived deformations generating $\epsilon=10^{-7}$ (r-mode and magnetic dipole radiation ignored). The right panel shows the variation of the braking index as the function of rotation frequency when the energy loss happens through the r-mode only with the amplitude $\alpha_{\rm S}=10^{-4}$. The solid yellow line represents one of the EOS from our sample (see Sec. IV.2 for the details), the blue dashed line indicates the DDME2 EOS and the black solid line indicates S271v6 EOS. The vertical magenta line indicates $f=716$ Hz line which is the maximum spin-frequency of a pulsar discovered to date.

### v2

@@ Abstract
Pulsars are rotating neutron stars that are observed to be slowing down, implying a loss of their rotational energy. There can be several different physical mechanisms involved in their spin-down process. The properties of fast-rotating pulsars depend on the nature of the neutron star matter, which can also affect the spin-down mechanisms. In this work, we examine three different physical phenomena contributing to the spin-down: magnetic dipole radiation, gravitational mass quadrupole radiation due to the “mountain” formation, gravitational mass current quadrupole radiation or the r-modes, and calculate the expressions for the braking indices due to all of them. We have also considered the implications of the uncertainties of the equation of the state of neutron star matter and rapid rotation on the braking indices corresponding to the aforementioned processes and their combinations. In all cases, the rapid rotation results in a departure of the braking index from the standard values when the rotational effects are ignored. If generated with a saturation amplitude within the range of $10^{-4}-10^{-1}$, the r-mode oscillations dominate the spin-down of millisecond pulsars. Moreover, we explore the braking index in the context of millisecond magnetars. We also study the effects of different choices of baryon mass on the braking indices.
@@ I Introduction
The majority of neutron stars (NS) are observed as pulsars, which are rapidly rotating and observed to spin down over time, indicating a loss of rotational kinetic energy. The spin-down rate is given as $\dot{\Omega}=-k\Omega^{n}$ Manchester et al. (1985), where $n$ is the braking index and $k$ is a positive constant. The braking index captures information about the mechanism of energy loss. Depending on the emission loss mechanism $k$ can be a function of radius, moment of inertia, magnetic field strength etc (discussed later in the text). In most cases, the pulsar’s rotational energy loss is attributed to the magnetic dipolar radiation (MDR) from the pulsar magnetosphere for which the braking index is 3 Gunn and Ostriker (1969). However, pulsars can lose energy via gravitational wave (GW) radiation if there is non-axisymmetric deformation or r-mode instability, for which the braking indices are 5 and 7, respectively (Riles, 2023). Therefore, it is natural to expect a rapidly rotating neutron star to lose energy via all three channels if the conditions are favourable. Apart from these three channels of energy loss, pulsars also lose energy due to the particle wind (Harding et al., 1999; Spitkovsky, 2006).
@@
The spin-period of pulsars spans a large range. The slowest pulsar has a rotation period of 76 s Caleb et al. (2022) to the most rapidly rotating being 1.39 ms Hessels et al. (2006). The total population is typically categorized as normal (or “slow”) pulsars and millisecond pulsars, with the demarcation around 16 ms Halder et al. (2023). The braking index has been measured for a few younger pulsars (through pulsar timing) and found to be less than 3 Livingstone et al. (2007); Weltevrede et al. (2011); Livingstone and Kaspi (2011); Lyne et al. (1993); Boyd et al. (1995); Roy et al. (2012); Espinoza et al. (2011). This is often attributed to the evolution of the magnetic field, modified magnetosphere or even the interaction of supernova fallback disk Chen and Ruderman (1993); Melatos (1997); Harding et al. (1999). A similar sub-3 braking index can be achieved in rapidly rotating pulsars, which is purely due to structural changes in the star due to rapid rotation. In such a scenario, the $k$ cannot be assumed independent of the spin frequency Hamil et al. (2015); Lan et al. (2021), which has been discussed in detail in Sec. IV.
@@
The effect of structural evolution is prominent over a spin frequency of 200 Hz Hamil et al. (2015). Currently, the ATNF Pulsar Catalogue  [^1] https://www.atnf.csiro.au/research/pulsar/psrcat/ Manchester et al. (2005) records 392 pulsars with spin-frequency more than 200 Hz, which corresponds to $\sim 11\%$ of the total pulsar population discovered till date. These rapidly rotating pulsars are important in probing the physics of dense nuclear matter. The physics of the nuclear interaction and the constituents of the matter determines the dense matter equation of state (EOS), which in turn governs the stellar structure and its response to rapid rotation. There are several ways found in the literature to model the NS EOS. One may use explicit nuclear potentials, or energy density functionals describing phenomenological nucleonic interactions Dutra et al. (2012, 2014); Oertel et al. (2017); Nandi et al. (2019); Sun et al. (2024). One may also employ more agnostic and semi-agnostic approaches which do not assume any interactions a priori Read et al. (2009); Lindblom (2010); Landry and Essick (2019); Margueron et al. (2018); Biswas et al. (2021). In this article, we have adopted a semi-agnostic approach to construct the dense matter equation of state following the formalism of Gandolfi et al. (2019), summarized in Sec. II, to study the frequency evolution of the stellar structure and the braking index.
@@
There are various types of constraints that can be applied to the EOS model. Astrophysical observations provide several constraints relevant to the high density part of the EOS. In particular, the radio observations of the most massive pulsar, PSR J0740+6620 Cromartie et al. (2019); Fonseca et al. (2021) provide the most stringent constraint on the NS EOS. The tidal deformability measurements from the multimessenger observations of the binary neutron star merger event GW170817 put additional constraints on the EOS Abbott et al. (2017a, b, 2019a). In recent years, the Neutron star Interior Composition Explorer (NICER) collaboration has reported a few simultaneous mass-radius measurements of PSR J0030+0451, PSR J0740+6620, and PSR J0437-4715, respectively, that also help us to constrain the Mass-radius plane and consequently the EOS parameter space Miller et al. (2019); Riley et al. (2019); Miller et al. (2021); Riley et al. (2021); Salmi et al. (2022); Vinciguerra et al. (2024); Choudhury et al. (2024). On the other hand, the latest constraint on the low-density part of the EOS comes from the advancements of the chiral effective field theory ($\chi$-EFT) calculations Tews et al. (2013); Hebeler et al. (2013); Lynn et al. (2016); Drischler et al. (2019); Huth et al. (2021).
@@
Figure 2: The brown lines indicate the 99% confidence interval of the constrained EOS corresponding to the posterior distributions of figure 1 and the blue lines indicate the EOS which falls within the 68% confidence interval of the constrained posterior distribution of the EOS parameters.
Figure 3: The M-R diagrams for all EOS within the 99% confidence interval of the posterior distribution from figure 2 have been shown in the brown coloured lines and the blue lines indicate the M-R relations corresponding to the 68% confidence interval of the posterior distribution of the constrained EOS parameters.
@@ II Description of neutron star matter
where, $P(m_{1},m_{2}|\mathbf{X})$ is the prior distribution for the component masses of the binary. For simplicity, we have chosen a uniform prior for $m_{1}$ and $m_{2}$. We use a Gaussian kernel density estimator to construct the GW likelihood from the discrete data. We have fixed the chirp mass to its median value because of the high precision of the measurement. Then, we construct binaries corresponding to that chirp mass by varying $m_{1}$ and determine the corresponding $m_{2}$. For each EOS, we compute the tidal deformabilities ($\Lambda_{1},\Lambda_{2}$) for the pairs of ($m_{1},m_{2}$), and find the probability using equation II.
@@
Finally, we sample the posterior using a nested sampling algorithm Skilling (2006), in the dynesty software package Speagle (2020). The posterior distributions of the EOS parameters for nuclear matter are shown in Fig. 1. We have also included the derived values for $E_{\rm sym}$ and $L$ at $\rho_{0}$. We have found $E_{\rm sym}(\rho_{0})=28.95^{+3.31}_{-2.17}$ MeV and $L=53.28^{+6.56}_{-5.16}$ MeV at 68% confidence interval (C.I.). These values are well within the agreed ranges found in literature Oertel et al. (2017); Margueron et al. (2018). This affirms the robustness of our EOS samples used to study the rotational properties of pulsars in this work. We have shown the corresponding EOSs and their mass-radius curves in Fig. 2 and 3, respectively. We can see from Fig. 3 that the EOSs generated in our exercise span a large range of radii for a $1.4M_{\odot}$ star. Therefore, these curves are able to represent the parameter space of neutron star matter. Note that, we have not used the data from NICER measurements in our analysis as we see that our 68% mass-radius curves in Fig. 3 are consistent with 68% of those sources.
@@ III Rotational effects on the structure of Neutron stars
The rapid rotation generates a large amount of centrifugal force, which leads to the departure from the spherical shape of the star that is characterized by the ratio of the polar to the equatorial radius $R_{p}/R_{e}$. The dependence of this ratio on frequency is shown in the bottom left panel of Fig. 4 for all EOS. Due to the deformation in the shape of the star, the definition of the radius becomes ambiguous in such a case, yet, in many scenarios, for estimating various physical quantities, the radius of the star is required as discussed in Sec. IV. Hence, in this work, we define a frequency-dependent radius by averaging over the circumferential radius of a rotating star given as
@@
The expression of circumferential radius $R(\theta,\Omega)=R_{e}[1-{\bar{\Omega}}^{2}(0.788-1.03x)\cos^{2}\theta]$ is taken from the previous work by AlGendy and Morsink (2014) and Suleimanov et al. (2020), where $\theta$ is the polar angle, $\bar{\Omega}=\Omega(R^{3}_{e}/GM)^{1/2}$ and $x=GM/c^{2}R_{e}$. The EOS and rotation frequency-dependent average radius of the star is shown in the lower right panel of Fig. 4, which can be seen to decrease strongly at higher frequencies ($\gtrsim 850$Hz). Also at the higher frequency, the $R_{p}$ can be $\sim 25\%$ smaller than the $R_{e}$, therefore our method of computing a unique value of a radius and attributing a spherical shape to the star may have some shortcomings even if it mathematically renders a valid solution. In the upper left and right panels of Fig. 4 we show the variation of the gravitational mass and moment of inertia as a function of the rotation frequency for all EOS. It is evident from Fig. 4 that all these physical quantities have significant frequency dependence and will contribute to the braking index of the NS which we will discuss in the next section. We have confined our studies to the rotational frequency above $\gtrsim 200$ Hz following Hamil et al. (2015), where the frequency dependence of the macroscopic quantities has been shown to be prominent above $\gtrsim 200$ Hz only.
@@
The high-frequency behaviour of the macroscopic quantities observed in Fig. 4 is controlled by the stiffness of the EOSs. At a given density, a stiffer EOS produces more pressure than a softer one and, therefore, can defy gravity more easily. Hence, if there is no rotation, the stiffer EOS will make a star of a larger radius and a lower central density if a certain number of baryons (equivalent to a fixed baryon mass) are put together. Consequently, the star with the stiffer EOS will be less bound, resulting in a higher gravitational mass. Note the gravitational binding energy is defined as $M_{b}-M$. This behaviour remains even if the stars rotate. Therefore, the upper curves in $M$-vs-$f$, $R$-vs-$f$, and $I$-vs-$f$ plots in Fig. 4 correspond to stiffer EOSs. With increasing frequency, the equatorial radii of all the stars (irrespective of the underlying EOS) grow, making them less bound and having higher gravitational mass. Furthermore, since at a given rotational frequency, the star with the stiffer EOS is less bound than the softer EOS, it becomes unstable at a lower frequency. In other words, the Kepler frequency is smaller in the case of stiffer EOSs. It explains the behaviour seen in the $M-f$ plot. The $R_{p}/R_{e}$-vs-$f$ curve captures the dependence of the stellar deformation on the rotation frequency. It is clear from the above discussion that for the stiffer EOSs, the equatorial radius increases much faster with the small change to the rotation frequency compared to the softer EOSs. Consequently, the ratio $R_{p}/R_{e}$ falls faster in the case of stiffer EOSs. Therefore, the left curves in the $R_{p}/R_{e}$-vs-$f$ plot belong to stiffer EOSs.
@@ IV.1 Calculation of braking index
$|\dot{E}_{\rm{EM}}|=\frac{2}{3c^{3}}\mu^{2}\Omega^{4}\sin^{2}\alpha\equiv\mathcal{E}_{EM}R^{6}\Omega^{4},$ (12)
where the magnetic moment $|\mu|^{2}=B_{p}^{2}R^{6}/4$, with the surface magnetic field strength $B_{p}$, stellar radius $R$, $\alpha$ is the angle of inclination between the rotation and the magnetic axis and $\mathcal{E}_{EM}=B_{p}^{2}\sin^{2}\alpha/6c^{3}$. The corresponding spin-down rate using Eq. 10 leads to
$\dot{\Omega}_{\rm MDR}=-\frac{2B_{p}^{2}R^{6}\sin^{2}\alpha}{3Ic^{3}}\Omega^{3}\Rightarrow-k_{\rm MDR}\Omega^{3},$ (13)
indicating the MDR leads to braking index 3. However, in Sec. III we have noticed at higher spin-frequency $R$ and $I$ depend on $f$ making $k_{\rm MDR}$ dependent on $f$ and hence the braking index. On the other hand, the $f$ dependence of $I$, will not allow us to arrive at the expression 13, as $dI/d\Omega\neq 0$. Hence, for a rapidly rotating star, the modification to the spin-down expression is essential.
@@
The expression of $n(\Omega)$ above shows the dependence of the braking index on the spin frequency, which is further dependent on the EOS through the frequency derivative of the macroscopic quantities such as $M$, $R$ and $I$. The term $n_{\rm EM}$ is the contribution of MDR loss, $n^{Q}_{\rm{GW}}$ is the contribution from GW radiation due to long-lived deformations, $n^{R}_{\rm{GW}}$ is the contribution from the emission loss via r-mode oscillations. The term $n_{I}$ depends purely on the moment of inertia and its higher-order derivatives, which has been previously derived in Glendenning et al. (1997). The quantities $\mathcal{E}_{EM},\epsilon$ and $\alpha_{\rm S}$ can be considered as switches which by setting to zero help to analyze the braking index in the absence of a particular channel of energy loss. For example, setting $\alpha_{\rm S}=0$ will help to understand the evolution of the braking index purely due to the electric dipolar emission loss and the GW emission loss due to finite $\epsilon$. In the limit of slow rotation where all the derivatives w.r.t $\Omega$ tend to zero, by setting $\epsilon$ and $\alpha_{\rm S}$ to zero we obtain $n=3$, by setting $\mathcal{E}_{EM}$ and $\alpha_{\rm S}=0$, we obtain $n=5$ and finally by setting $\mathcal{E}_{EM}=\epsilon=0$, we obtain $n=7$, which are in agreement with the limiting cases known in the literature.
[…]